OpenStudy (anonymous):

Write the standard form an equation of the line passing through the given point with the given slope. m=7/4; (3,-1)

OpenStudy (anonymous):

ok standard form means Ax+By=C

OpenStudy (anonymous):

we have y--1=7/4(x-3) --->y+1=7/4(x-3)

OpenStudy (anonymous):

A has to be positive integer

OpenStudy (anonymous):

y-7/4x=-25/4

OpenStudy (anonymous):

-7/4 x+y=-25/4

OpenStudy (anonymous):

ok multiply everything by -4...to make coefficient of A positive and all others an integer

OpenStudy (anonymous):

7x-4y=25

OpenStudy (anonymous):

thats the answer @lilli.tharp